OpenStudy (anonymous):

Plz Help Me... I'll give medal.. Hallie is trying to win the grand prize on a game show. Should she try her luck by spinning a wheel with 6 equal sections labeled from 1 to 6 and hope she gets a 5, or should she roll two number cubes and hope she gets the same number on both cubes? Explain..

OpenStudy (anonymous):

I'm really not good at these..

OpenStudy (anonymous):

Either method is equally likely. They both will produce a success at: Probability = 1/6 The spin will give success at 1 out of 6 The dice roll will give success at 6 out of 36 -> 1 out of 6 They are equal as far as probability goes.

OpenStudy (im_an_angel):

@nono266 is exactly right :)

OpenStudy (mathstudent55):

Not quite.

OpenStudy (mathstudent55):

You need to analyze each situation.

OpenStudy (anonymous):

So Mathstudent55 what do you think?

OpenStudy (anonymous):

oh @mathstudent55 haha thanks @im_an_angel

OpenStudy (tkhunny):

First, what numbers are on the number cubes? Are there any matches? I still am amazed that some are offended by a reference to "Standard six-sided dice".

OpenStudy (anonymous):

I'm really not good at Probability

OpenStudy (mathstudent55):

If you spin a wheel with 6 equally likely results, then the probability of any of them is 1/6.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

No. I stand corrected. You were right in the first place. With the dice it's the same thing. There are 36 total outcomes, and 6 of them are doubles, so the probability is indeed 6/36 = 1/6.

OpenStudy (anonymous):

oh. ok

OpenStudy (anonymous):

How should I explain it?

OpenStudy (anonymous):

Are there eny awnsers

OpenStudy (mathstudent55):

@Madi_Lee Do you understand why with the wheel the probability is 1/6?

OpenStudy (anonymous):

because of 1 out of 6?

OpenStudy (mathstudent55):

Yes.

OpenStudy (mathstudent55):

With the dice, you spin both dice. There are 6 possible outcomes of the first die, and there are 6 possible outcomes of the second die. The total number of outcomes is 6 * 6 = 36 (For each number you get in one die, you can get 6 numbers on the second one. Multiply them, and you get 36 total possible outcomes.)

OpenStudy (anonymous):

Oh.. Ok

OpenStudy (mathstudent55):

Now you need to see how many outcomes you are interested in. The doubles spins are: 1,1; 2,2; 3,3; 4,4; 5,5; 6,6 As you can see, there are 6 possible doubles spins. 6 out of 36 is the same as 1 out of 6.

OpenStudy (anonymous):

Oh. That's it?

OpenStudy (anonymous):

Thank you

OpenStudy (mathstudent55):

Yes. That's it.